OpenStudy (anonymous):

Find the smallest positive N such that("==" means congruent). N==6(mod12) N==6(mod18) N==6(mod24) N==6(mod30) N==6(mod60)

OpenStudy (kittiwitti1):

http://www.math.cornell.edu/~putnam/modular.pdf

OpenStudy (anonymous):

Hi, my understanding is that if I tke last two congruences, I get: 30a+6=60b+6, which simplifies to 30a + 60 b so when I get Mod30 by guess and check I get 0 mod30 Is my Ans correct?

ganeshie8 (ganeshie8):

simply take LCM of moduli and add 6

OpenStudy (anonymous):

Well... We can say that: $$ a \equiv b \mod n \implies a-b = 0 \mod n $$Which means that \(n\) divides \(a-b\) So those turn into: $$ N \equiv 6 \mod 12 \implies N - 6 \equiv 0 \mod 12 \\ N \equiv 6 \mod 18 \implies N - 6 \equiv 0 \mod 18 \\ N \equiv 6 \mod 24 \implies N - 6 \equiv 0 \mod 24 \\ N \equiv 6 \mod 30 \implies N - 6 \equiv 0 \mod 30 \\ N \equiv 6 \mod 60 \implies N - 6 \equiv 0 \mod 60 $$Which means that \(N-6\) has to be a number that 12, 18, 24, 30 and 60 divide. So to find it our work is similar to finding the `Least common multiple` of numbers: http://en.wikipedia.org/wiki/Least_common_multiple And we can do it by prime factorization (as described there as well). $$ 12 = 3 \cdot 2 \cdot 2 \qquad 18 = 3 \cdot 3 \cdot 2 \\ 24 = 3 \cdot 2 \cdot 2 \cdot 2 \quad 30 = 3 \cdot 5 \cdot 2 \\ 60 = 3 \cdot 5 \cdot 2 \cdot 2 \\ $$So we get that \(N-6\) has to be multiply of 360 in order to divide all these numbers: $$ N-6 = k\cdot(2 \cdot 2 \cdot 2 \cdot 3 \cdot 3 \cdot 5) = k \cdot 360\\ N = k \cdot 360 + 6 $$By setting \(k=0\) we get \(N=6\) which is the smallest positive number that makes it work.. but that is the trivial solution.. In case you need something more interesting you can try \(k=1 \implies N=366\)

OpenStudy (mathmath333):

\(\large \begin{aligned} \color{black}{ x\equiv 6\pmod {360} }\end{aligned}\)

OpenStudy (anonymous):

but 366 is not the correct answer

OpenStudy (mathmath333):

oh wait it should be simply \(\Large 6\) then.

ganeshie8 (ganeshie8):

366 is the smallest positive integer

OpenStudy (anonymous):

6! I got it! in the second try, lost 2 points... thanks!

OpenStudy (mathmath333):

but 6 is also the remainder when divided by such numbers , so i think ir should be 6

ganeshie8 (ganeshie8):

ahh k=0, 360k + 6 = 6 will do !

OpenStudy (mathmath333):

so actually it was an observation question

OpenStudy (anonymous):

I know, but the funny thing is that for N=6 no work should be done really. it's the instant answer because 6 is smaller than all of the other numbers 12, 18, 24, 30, 60... so its modulo is the same for all

OpenStudy (anonymous):

Because the moduli are not relatively prime, we do not have general algebraic tools to approach this problem. However, we can see by inspection that 6 clearly satisfies all five of the congruences and that no positive integer less than 6 could satisfy them

ganeshie8 (ganeshie8):

a | 0 any number divides 0.. yeah its more of an observation q

OpenStudy (mathmath333):

"the smallest positive N"

OpenStudy (anonymous):

thank you! mathmath333 and ganeshie8 and pitamar!